OpenStudy (anonymous):

which line is perpendicular to 8x+3y=8?

OpenStudy (anonymous):

a. -3x-8y=8 b.-3x+8y=8 c.-8x+3y=8 d. 8x-3y=8

OpenStudy (anonymous):

well since perpendicular lines have the opposite reciprocal slopes we have to change 8x+3y=8 to slope intercept form. 3y=-8x+8 y=-8/3x+8/3 the slope of the line perpendicular to it has to be the opposite reciprocal. so the opposite of -8/3 is 8/3 and the reciprocal is 3/8. so the slope of the line perpendicular to it is 3/8 and b is the only answer with the slope of 3/8. Answer: B

OpenStudy (anonymous):

@lacrosseplayer22 thank you so much!